a varies directly as b; and a = 4 when b = 24
a) write an equation that relates a and b
b/a=24/4=6
b)find a when b=63
63/4=15.75
c) find b when a=7
24/7= 3.4286

Answer by stanbon(75887)

a varies directly as b; and a = 4 when b = 24
a) write an equation that relates a and b
a = k*b
Solve for "k"
4 = k*24
k = 1/6
Equation:
a = (1/6)b
---------------------
b)find a when b=63
a = (1/6)*63 = 63/6 = 10 1/2
-----------------------------
c) find b when a=7
7 = (1/6)b
b = 6*7 = 42
--------------------
Cheers,
Stan H.
----------------